Thursday, 9 am MT, 11 am ET, 11 December 2014

Logistics

CV-331/AOC-317/GB-bsm, 331 Hub 192.33.117.12##6519 Audio 434-817-6286

Agenda

  • 15B
    • CfP status (All)
    • Timeline (Dana)
    • PST Testing (Dana, Jen, Rick)
    • External Tools
      • GOST (Gustaaf)
      • VLA Exposure Calc. (Mark)
      • GBT Sensitivity Calc. (Toney)
    • PHT (Gustaaf/Barry)
    • Documentation (All)
    • SRP Recruitment (Mark/Dave)

  • AOB

Minutes

Attending: BarryClark, RickLively, JenPlank, HeidiMedlin, MarkClaussen, MeriStanley, ClaireChandler, RickAn, GustaafVanMoorsel, LorantSjouwerman, DanaBalser, ToneyMInter

  • 15B
    • CfP status. The telescope sections are almost complete. Still working on Intro and News sections.
    • Timeline. ToneyMInter would like to have two weeks (instead of one week) between the time we finalize the SRPs and when we release info to the TAC. This extra time is required to accommodate the uncertainty in dealing with several potential new GBT sponsors. The issue is in generating a schedule which will depend on the details of these agreements. This may only be an issue for 15B but this was not clear. Everyone agreed that it would be difficult to make this up anywhere else in the process. SRP and TAC members have been asking about the schedule so we need to converge soon.
    • PST Testing. No progress on general testing for VLA/VLBA (MarkClaussen) or GBT (DavidFrayer). All items in the Plan of Record have been completed and had some testing. Everyone should try the new PDF generation for issues and report any problems to DanaBalser. We still need to perform stress tests on the PDF generation and some additional tests on the semester date/time deadline.
    • External Tools
      • GOST. LorantSjouwerman reported that there are only a few minor changes and bug fixes. No testing has been done yet. It does not appear that there will be a java update but we should confirm that GOST works on all NRAO systems.
      • VLA Exposure Calc. MarkClaussen reported that they are working on some bug fixes but no major issues.
      • GBT Sensitivity Calc. ToneyMinter reported no changes for the GBT.
    • PHT. BarryClark and GustaafVanMoorsel will review the outstanding JIRA tickets and form a new Plan of Record.
    • Documentation. Only minor updates are required for proposal documentation.
    • SRP Recruitment. MarkClaussen reported that there are about 20 slots to fill. He solicited the scientific staff for recommendation and has heard back from several people.

  • AOB
    • GBT JPL Radar Back-end. This back-end does not need approval by the PI. There was some discussion about how the various approval checkboxes were behaving in the PST. DanaBalser and RickLively will resolve these.

-- DanaBalser - 2014-12-10
Topic revision: r5 - 2014-12-18, DanaBalser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback